Solve log(2x+3) = log(4x) + 2 for x.
Answer by stanbon(48535)

Subtract "log(4x)" from each side to get
log(2x+3)-log(4x)=2
Apply the Quotient law of logarithms to get
log[(2x+3)/4x]=2
The corresponding exponental equation is
[(2x+3)/4x]=10^2 = 100
Multiply both sides by 4x to get the following
2x+3=400x
398x=3
x= 3/398
Cheers,
Stan H.